# Amplifying hybrid entangled states and superpositions of coherent states

### v1

@@ Abstract
We compare two amplification schemes, photon addition and then subtraction ($\hat{a}\hat{a}^{\dagger}$) and successive photon addition ($\hat{a}^{\dagger}{}^{2}$), applied to hybrid entangled states (HESs) and superpositions of coherent states (SCSs). We show that the amplification schemes’ fidelity and gain for HESs are the same as those of coherent states. On the other hand, SCSs show quite nontrivial behaviors by the amplification schemes, depending on the amplitudes of coherent states, number of coherent-state components, and relative phases between the components. This implies that appropriate amplification schemes for SCSs should be chosen depending on the tasks and specific forms of the states. To investigate the quality of amplified states, we calculate the quantum Fisher information, a measure of quantum phase estimation. In terms of the quantum Fisher information, the $\hat{a}\hat{a}^{\dagger}$ scheme tends to show better performance for relatively small amplitudes while the $\hat{a}^{\dagger}{}^{2}$ scheme is better in larger amplitude regime. The performance of the two schemes becomes indistinguishable as the amplitude grows sufficiently large.
@@
Hybrid entangled states (HESs) Jeong14 ; Morin14 and superpositions of coherent states (SCSs) Yurke86 ; Schleich91 in free-traveling fields are well-known nonclassical states. A HES is an entangled state between a discrete-variable (DV) system such as a photon number state and a continuous-variable (CV) system such as a coherent state. This integrated perspective toward the traditional dichotomy of CV and DV quantum information brings new advantages that could not have been obtained in each area individually. Apart from its theoretical versatility, interplay and conversion between CV and DV systems Jeong23 have been experimentally realized Ulanov17 ; Sychev18 ; Darras23 indicating generation and processing of hybrid entanglement are well within our reach. HESs are also useful for fundamental studies on quantum physics such as loss-tolerant verification of Bell-inequality violations Kwon13 , quantum teleportation Park12 ; Lee13 ; Jeong16 ; Kim16 ; Ulanov17 ; Sychev18 , quantum computation Lee13 ; Omkar19 , quantum communication Bose23 ; Lim16 ; JK2001 and steering verification Cavailles18 ; Huang19 . Generation schemes for HESs were theoretically proposed Gerry99 ; Nemoto04 ; Jeong05a ; Anderson13 ; Jeong14 ; Morin14 and experimentally implemented Jeong14 ; Morin14 . SCSs are superpositions of coherent states with different phases, and a typical example is a superposition of two coherent states with opposite phases. They are useful for many quantum information tasks such as quantum computation Cochrane99 ; Jeong02 ; Ralph03 ; Lund08 ; Myers11 ; Kim15 , quantum teleportation vanEnk01 ; Jeong01 ; Neergaard-Nielsen13 , precision measurements Gerry01 ; Gerry02 ; Ralph02 ; Munro02 ; Campos03 ; Joo11 ; Hirota11 ; Joo12 ; Zhang13 , non-Gaussian state generation  Vasconcelos10 ; Weigand18 ; Shunya24 and error suppression Li17 ; Hastrup22 . They have also been used to investigate violations of local realism Wilson02 ; Park15 and nonlocal realism Paternostro10 ; Lee11 . SCSs in free-traveling fields have been implemented approximately using various methods Ourjoumtsev06 ; Neergaard-Nielsen06 ; Ourjoumtsev07 ; Wakui07 ; Ourjoumtsev09 ; Jeong05 In these regards, obtaining sufficiently large amplitudes of continuous variables in HESs or SCSs is required to accomplish various quantum information processing tasks.
@@
One way to obtain HESs or SCSs of large amplitudes is to amplify the states Ralph03 ; Lund04 ; Neergaard-Nielsen13 ; Lag13 ; Lvovsky2017 ; Oh18 or purify them JK2001 ; Sheng13 . Although a deterministic noiseless linear amplification for a bosonic system is forbidden by quantum mechanics Caves88 , there have been studies on nondeterministic amplification of coherent states Parigi07 ; Zavatta11 ; Park16 . In this paper, we adopt the photon addition and then subtraction scheme Parigi07 ; Zavatta11 ($\hat{a}\hat{a}^{\dagger}$) and the successive photon addition scheme Park16 ($\hat{a}^{\dagger}{}^{2}$). We compare the fidelity and gain of the two schemes and take quantum Fisher information for phase estimation  Helstrom76 ; Braunstein94 ; Barndorff-Nielsen00 ; Paris09 ; Petz10 as an indicator for the performance of amplification. This could be a practical measure in estimating the amplification performance because various nonclassical states of light fields are utilized for quantum phase estimation Usuga10 ; Genoni11 ; Berni15 ; Izumi16 . In fact, in a previous work Park16 , the quantum Fisher information is employed as a measure for amplification schemes using $\hat{a}\hat{a}^{\dagger}$ and $\hat{a}^{\dagger}{}^{2}$ for $2$-dimensional SCSs. Extending SCSs to multi-dimensional cases would be consistent with employing the same measure to show the performance of amplified states. However, we do not adopt the equivalent-input-noise (EIN) Grangier92 as a measure of noiseless amplification because the expectation value of a quadrature operator, $\langle x_{\lambda}\rangle$, is zero for all $\lambda$ for hybrid states and SCSs.
@@
Consequently, we show that the photon addition and subtraction scheme gives higher fidelity while the $(\hat{a}^{\dagger})^{2}$ scheme gives higher gain. Performance of optimal quantum phase estimation measured by quantum Fisher information shows that photon addition and then subtraction scheme is better for relatively small amplitude, where its exact value differs by quantum states, $(\hat{a}^{\dagger})^{2}$ scheme is better for larger amplitude, and the difference of their performance vanishes as amplitude increases.
@@
In Sec. II, we begin with preliminaries of HESs and SCSs. In Sec. III, we investigate and compare the two amplification schemes applied to hybrid states, and in Sec. IV, we investigate and compare those schemes applied to SCSs. In Sec. V, we conclude our paper with assessments of the two amplification schemes.
@@ II.1 Hybrid entangled states
Hybrid entangled states(HES) are bosonic states consisting of entanglement between CV states and DV states. we define a qubit basis of HESs here as
@@
where |$\ket{n,\alpha}=\ket{n}\otimes\ket{\alpha}$, a product of a number state with photon number $n$ and a coherent state with amplitude $\alpha$($\in\mathbb{R}^{+}$ for convenience). For a 2D HES system, The DV part can also be polarization states |$\ket{H}$ and |$\ket{V}$. Although coherent states |±$|\pm\alpha\rangle$ are not orthogonal to each other, orthogonality of number states provides independence of normalization factor with respect to amplitude $\alpha$.
HES qubits can be generalized to represent qudits using d-fold complex roots of 1, $\omega_{d}:=e^{\frac{2\pi i}{d}}$.
@@
Here $k$ represents the index for the d-dimensional qudit. Overlap between arbitrary basis states can be calculated as
@@
$=\frac{1}{d}\sum_{n}\omega_{d}^{(k-l)n}=\delta_{k,l}.$ (3)
which indeed shows orthonormality. Also, the photon number statistics of HES are the same with coherent states with the same amplitude. We do not explicitly prove this since the statement can be directly derived from Proposition 1 in Sec. III.
Another important feature of the HES qudits is that photon addition/subtraction maps them to other HES qudits. For a $d$-dimensional hybrid state |$\ket{\mathcal{H}^{k}_{\alpha,d}}$, $m$ $(\hat{a}^{\dagger})^{2}$ maps |$\ket{\mathcal{H}^{k}_{\alpha,d}}$ to |$\ket{\mathcal{H}^{l}_{\beta,d}}$ for some different amplitude $\beta$ where $l\equiv k+m(\mathrm{mod}\ d)$.
@@
$=\frac{1}{\mathcal{N}}\delta_{l,k+m;d}\,\beta^{m}\mathrm{Exp}[-\frac{1}{2}(\alpha-\beta)^{2}].$ (4)
where $\delta_{m,k;d}:=\delta_{m,k+d\cdot j}$ with $j\in\mathbb{Z}$. That is, $\delta_{m,k;d}=1$ if and only if $m\equiv k(\mathrm{mod}\ d)$. Also, $\mathcal{N}=\mathcal{N}(\alpha)=\sqrt{\sum_{k=0}^{n}\frac{(n!)^{2}}{k!((n-k)!)^{2}}\alpha^{2(n-k)}}$ is a normalization factor from photon additions. The normalization factor $\mathcal{N}\approx\alpha^{n}$ in the asymptotic limit of $\alpha\gg 1$. In such a limit the optimal choice of $\beta$ yielding maximal overlap becomes $\alpha$, and in that case, fidelity is strictly 1. Even if we consider a more realistic regime with smaller amplitudes($\alpha<9$), we can achieve $F>0.99$ with well-chosen $\beta$ up to $\approx$12 successive photon addition.
@@ II.2 Superposition of coherent states
Superposition of coherent states(SCS), as the name suggests, is a linear combination of classical coherent states with distinct phases. Typical SCS qubit basis, also known as even/odd cat states are given by
@@
It is worth noting that individual coherent components are not orthogonal, making the normalization factor of SCS depend on the coherent state amplitude $\alpha$. Moreover, orthogonality between |$\ket{\mathcal{C}^{0}_{\alpha}}$ and |$\ket{\mathcal{C}^{1}_{\alpha}}$ is strictly guaranteed by relative phase between their components regardless of their nonorthogonality. This can be explicitly proven by attaining the number basis expansion of 2D SCSs.
@@
$N_{k,\alpha,d}=\frac{1}{\sqrt{d\sum_{n}\omega_{d}^{-kn}\mathrm{Exp}[-\alpha^{2}(1-\omega_{d}^{n})]}}.$ (8)
@@
$\begin{split}\bra{m}\ket{\mathcal{C}^{k}_{\alpha,d}}&=N_{k,\alpha,d}\sum_{n}\omega_{d}^{-kn}\bra{m}\ket{\alpha\omega_{d}^{n}}\\ &=N_{k,\alpha,d}\sum_{n}\omega_{d}^{-kn}\frac{(\alpha\omega_{d}^{n})^{m}}{\sqrt{m!}}\\ &=\frac{\alpha^{m}N_{k,\alpha,d}}{\sqrt{m!}}\sum_{n}\omega_{d}^{n(m-k)}=\frac{\alpha^{m}N_{k,\alpha,d}}{\sqrt{m!}}\delta_{m,k;d}\end{split}$ (9)
@@
III Amplification of Hybrid State
In this section, we compare the two amplification schemes applied to hybrid states. The first is addition-subtraction, namely $\hat{a}\hat{a}^{\dagger}$-scheme, while the second comprises successive photon addition, namely $a^{\dagger}{}^{2}$-scheme. Deterministic implementation of both schemes is fundamentally forbidden since they are non-unitary. Thus, quantum channels enabling the probabilistic application of those operations are designed to realize those schemes.
@@ III Amplification of Hybrid State
|$\begin{split}&\ket{\mathcal{H}^{k,\hat{a}\hat{a}^{\dagger}}_{\alpha,d}}=N_{hyb}^{\hat{a}\hat{a}^{\dagger}}(\alpha,d)\,(I\otimes\hat{a}{\hat{a}^{\dagger}})\frac{1}{\sqrt{d}}\sum_{n}\omega_{d}^{-kn}\ket{n,\alpha\omega_{d}^{n}}\\ &\ket{\mathcal{H}^{k,(\hat{a}^{\dagger})^{2}}_{\alpha,d}}=N_{hyb}^{\hat{a}^{\dagger}{}^{2}}(\alpha,d)\,(I\otimes(\hat{a}^{\dagger})^{2})\frac{1}{\sqrt{d}}\sum_{n}\omega_{d}^{-kn}\ket{n,\alpha\omega_{d}^{n}}\end{split}$ (10)
where $N_{hyb}^{\hat{a}\hat{a}^{\dagger}}(\alpha,d),N_{hyb}^{(\hat{a}^{\dagger})^{2}}(\alpha,d)$ normalization factors required for the consequent states to be normalized. Note that we only operate on CV parts.
$N_{hyb}^{\hat{a},\hat{a}^{\dagger}}(\alpha,d)=\frac{1}{\sqrt{\alpha^{4}+3\alpha^{2}+1}}$ (11)
$N_{hyb}^{\hat{a}\dagger^{2}}(\alpha,d)=\frac{1}{\sqrt{\alpha^{4}+4\alpha^{2}+2}}$ (12)
@@
$\langle P(\hat{a},\hat{a}^{\dagger})\rangle_{\mathcal{H}^{k}_{\alpha,d}}=\langle P(\hat{a},\hat{a}^{\dagger})\rangle_{\alpha}$ (13)
Trace preserving property of quantum channels and Proposition 1 gives an expression for normalization factor $N_{hyb}^{P(\hat{a},{\hat{a}^{\dagger}})}(\alpha,d)$:
$N_{hyb}^{P(\hat{a},{\hat{a}^{\dagger}})}(\alpha,d)=\frac{1}{\sqrt{\bra{\alpha}P(\hat{a},{\hat{a}^{\dagger}})^{\dagger}P(\hat{a},{\hat{a}^{\dagger}})\ket{\alpha}}},$ (14)
@@
we take fidelity and gain as our figures of merit to compare the two schemes($\hat{a}\hat{a}^{\dagger},(\hat{a}^{\dagger})^{2}$). Fidelity can be considered as a measure of how close we can get to our target state, while gain quantifies our capability to increase the amplitude of the component states. Let us denote the amplitude of the target state as $g\alpha$ where $g$ is the gain value. For the photon addition and then subtraction scheme, the fidelity is analytically obtained as
$F^{\hat{a}\hat{a}^{\dagger}}=\frac{g^{2}\alpha^{4}+2g\alpha^{2}+1}{\alpha^{4}+3\alpha^{2}+1}\mathrm{Exp}[-\alpha^{2}(g-1)^{2}],$ (15)
@@
$F^{\hat{a}^{\dagger}{}^{2}}=\frac{g^{4}\alpha^{4}}{\alpha^{4}+4\alpha^{2}+2}\mathrm{Exp}[-\alpha^{2}(g-1)^{2}].$ (16)
Figure 1: (a) Fidelity graphs of the two schemes, $\hat{a}\hat{a}^{\dagger}$ and $\hat{a}^{\dagger}{}^{2}$. Blue line denotes $F_{hyb}^{\hat{a}\hat{a}^{\dagger}}$ while orange line denotes $F_{hyb}^{\hat{a}^{\dagger}{}^{2}}$. (b) Gain graphs of the two schemes. Blue line denotes $G_{hyb}^{\hat{a}\hat{a}^{\dagger}}$ while orange line denotes $G_{hyb}^{\hat{a}^{\dagger}{}^{2}}$. (c) Ratio of quantum Fisher information of amplified states over those of hybrid states, $\mathcal{F}_{N}$. Blue line denotes $\mathcal{F}_{hyb}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{hyb}$ while orange line denotes $\mathcal{F}_{hyb}^{\hat{a}^{\dagger}{}^{2}}/\mathcal{F}_{hyb}$. (d) Ratio of quantum Fisher information of the two schemes, $\mathcal{R}=\mathcal{F}_{hyb}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{hyb}^{\hat{a}^{\dagger}{}^{2}}$.
The fidelity graphs of the two schemes are shown in Figure 1. (a). One can see that the $\hat{a}\hat{a}^{\dagger}$ scheme yields higher fidelity than $(\hat{a}^{\dagger})^{2}$ scheme. We first note that the fidelity in Eq. (15) and Eq. (16) are the same for the fidelity of coherent states due to Proposition 1. Explicit calculation of fidelities in Appendix A indeed has the form of Eq. (12). The behavior of fidelity between the target coherent state with greater amplitude and the coherent state amplified with either $\hat{a}\hat{a}^{\dagger}$ scheme or $(\hat{a}^{\dagger})^{2}$ scheme are well known Park16 .
@@
$\langle\mathcal{H}_{\beta,d}^{k+l}|Q(\hat{a},\hat{a}^{\dagger};l)|\mathcal{H}_{\alpha,d}^{k}\rangle=\langle\beta|Q(\hat{a},\hat{a}^{\dagger};l)|\alpha\rangle$ (18)
Unlike Proposition 1, this agreement between fidelity of hybrid state amplification and coherent state amplification does not solely originate from the orthogonality of discrete variables. We both need discrete variables |$\ket{n}$ which act as labels of continuous variables |$\ket{\alpha\omega_{d}^{n}}$ and equal phase factor $\omega$ which constitutes both qudit number $k$ by relative phase $\omega^{-kn}$ and phase of coherent states |$\ket{\alpha\omega^{n}}$ to cancel out relative phase.
Now let us obtain the gain value defined as the ratio between the amplitudes of the target state to the initial state giving the maximum fidelity. This can be obtained by solving the following equation

### v2

@@ Abstract
We compare two amplification schemes, photon addition and then subtraction ($\hat{a}\hat{a}^{\dagger}$) and successive photon addition ($\hat{a}^{\dagger}{}^{2}$), applied to hybrid entangled states (HESs) and superpositions of coherent states (SCSs). We show that the fidelity and gain of the amplification schemes for HESs are the same as those of coherent states. However, SCSs show quite non-trivial behavior by the amplification schemes, depending on the amplitudes of coherent states, the number of coherent-state components, and the relative phases between the components. This implies that appropriate amplification schemes for SCSs should be chosen depending on the tasks and specific forms of the states. To investigate the quality of amplified states, we calculate the quantum Fisher information, a measure of quantum phase estimation. In terms of the quantum Fisher information, the $\hat{a}\hat{a}^{\dagger}$ scheme tends to show better performance for relatively small amplitudes ($\alpha\lesssim 0.9$), while the $\hat{a}^{\dagger}{}^{2}$ scheme is better in the larger amplitude regime. The performance of the two schemes becomes indistinguishable as the amplitude grows sufficiently large.
@@
Hybrid entangled states (HESs) Jeong14 ; Morin14 and superpositions of coherent states (SCSs) Yurke86 ; Schleich91 in free-traveling fields are well-known nonclassical states. A HES is an entangled state between a discrete-variable (DV) system such as a photon number state and a continuous-variable (CV) system such as a coherent state. This integrated perspective toward the traditional dichotomy of CV and DV quantum information brings new advantages that could not have been obtained in each area individually. Apart from its theoretical versatility, interplay, and conversion between CV and DV systems Jeong23 have been experimentally realized Ulanov17 ; Sychev18 ; Darras23 indicating generation and processing of hybrid entanglement are well within our reach. HESs are also useful for fundamental studies on quantum physics such as loss-tolerant verification of Bell-inequality violations Kwon13 , quantum teleportation Park12 ; Lee13 ; Jeong16 ; Kim16 ; Ulanov17 ; Sychev18 , quantum computation Lee13 ; Omkar19 , quantum communication Bose23 ; Lim16 ; JK2001 and steering verification Cavailles18 ; Huang19 . Generation schemes for HESs were theoretically proposed Gerry99 ; Nemoto04 ; Jeong05a ; Anderson13 ; Jeong14 ; Morin14 and experimentally implemented Jeong14 ; Morin14 . SCSs are superpositions of coherent states with different phases, and a typical example is a superposition of two coherent states with opposite phases. They are useful for many quantum information tasks such as quantum computation Cochrane99 ; Jeong02 ; Ralph03 ; Lund08 ; Myers11 ; Kim15 , quantum teleportation vanEnk01 ; Jeong01 ; Neergaard-Nielsen13 , precision measurements Gerry01 ; Gerry02 ; Ralph02 ; Munro02 ; Campos03 ; Joo11 ; Hirota11 ; Joo12 ; Zhang13 , non-Gaussian state generation  Vasconcelos10 ; Weigand18 ; Shunya24 and error suppression Li17 ; Hastrup22 . They have also been used to investigate violations of local realism Wilson02 ; Park15 and nonlocal realism Paternostro10 ; Lee11 . SCSs in free-traveling fields have been implemented approximately using various methods Ourjoumtsev06 ; Neergaard-Nielsen06 ; Ourjoumtsev07 ; Wakui07 ; Ourjoumtsev09 ; Jeong05 In these regards, obtaining sufficiently large amplitudes of continuous variables in HESs or SCSs is required to accomplish various quantum information processing tasks.
@@
One way to obtain HESs or SCSs of large amplitudes is to amplify the states Ralph03 ; Lund04 ; Neergaard-Nielsen13 ; Lag13 ; Lvovsky2017 ; Oh18 or purify them JK2001 ; Sheng13 . Although deterministic and noiseless linear amplification for a bosonic system is forbidden by quantum mechanics Caves88 , there have been studies on non-deterministic amplification of coherent states Parigi07 ; Fadrny24 ; Zavatta11 ; Park16 . In this paper, we adopt the photon addition and then subtraction scheme Parigi07 ; Zavatta11 ($\hat{a}\hat{a}^{\dagger}$) and the successive photon addition scheme Park16 ; Fadrny24 ($\hat{a}^{\dagger}{}^{2}$). We compare the fidelity and gain of the two schemes and take the quantum Fisher information for phase estimation  Helstrom76 ; Braunstein94 ; Barndorff-Nielsen00 ; Paris09 ; Petz10 as an indicator of amplification performance. This could be a practical measure to estimate the amplification performance because various non-classical states of light fields are utilized to estimate the quantum phase Usuga10 ; Genoni11 ; Berni15 ; Izumi16 . In fact, in a previous work Park16 , the quantum Fisher information is used as a measure for amplification schemes using $\hat{a}\hat{a}^{\dagger}$ and $\hat{a}^{\dagger}{}^{2}$ for $2$-dimensional SCSs. Extending SCSs to multidimensional cases would be consistent with employing the same measure to show the performance of amplified states. However, we do not adopt the equivalent-input-noise (EIN) Grangier92 as a measure of noiseless amplification because the expectation value of a quadrature operator, $\langle x_{\lambda}\rangle$, is zero for all $\lambda$ for HESs and SCSs.
@@
Consequently, we show that the $\hat{a}\hat{a}^{\dagger}$ scheme gives higher fidelity while the $\hat{a}^{\dagger}{}^{2}$ scheme gives higher gain. Performance of optimal quantum phase estimation measured by the quantum Fisher information shows that the $\hat{a}\hat{a}^{\dagger}$ scheme is better for relatively small amplitudes, where the exact threshold value differs with respect to quantum states. The $\hat{a}^{\dagger}{}^{2}$ scheme is better for larger amplitudes, and the difference between their performances asymptotically vanishes as the input-state amplitude increases.
@@
In Sec. II, we begin with preliminaries of HESs and SCSs. In Sec. III, we investigate and compare the two amplification schemes applied to HESs, and in Sec. IV, we investigate and compare those schemes applied to SCSs. Analyses of the success probabilities of amplification schemes acting on HESs and SCSs are provided in Sec. V. In Sec. VI, we conclude our paper by comparing the two amplification schemes.
@@ II.1 Hybrid entangled states
The HESs are bosonic states consisting of entanglement between CV states and DV states. We define a qubit basis of HESs here as
@@
where |$\ket{n,\alpha}=\ket{n}\otimes\ket{\alpha}$, a product of a number state with photon number $n$ and a coherent state with the amplitude $\alpha$($\in\mathbb{R}^{+}$ for convenience). For a 2D HES system, The DV part can also be polarization states |$\ket{H}$ and |$\ket{V}$. Although coherent states |±$|\pm\alpha\rangle$ are not orthogonal to each other, orthogonality of number states provides independence of normalization factor with respect to the amplitude $\alpha$.
HES qubits can be generalized to represent qudits using d-fold complex roots of 1, $\omega_{d}:=e^{\frac{2\pi i}{d}}$,
@@
where $0\leq k<d$ and $k$ is a non-negative integer index for the $d$-dimensional qudit basis states |$\ket{\mathcal{H}^{k}_{\alpha,d}}$. The overlap between arbitrary basis states can be calculated as
@@
$=\frac{1}{d}\sum_{n}\omega_{d}^{(k-l)n}=\delta_{k,l},$ (3)
which indeed shows orthonormality. Also, the photon number statistics of HESs are the same with coherent states with the same amplitude. We do not explicitly prove this since the statement can be directly derived from Proposition 1 in Sec. III.
Another important feature of the HES qudits is that photon addition or subtraction maps them to other HES qudits. For a $d$-dimensional HES |$\ket{\mathcal{H}^{k}_{\alpha,d}}$, $m$ successive photon addition maps |$\ket{\mathcal{H}^{k}_{\alpha,d}}$ to |$\ket{\mathcal{H}^{l}_{\beta,d}}$ for some different amplitude $\beta$ where $l\equiv k+m(\mathrm{mod}\ d)$.
@@
$=\frac{1}{\mathcal{N}}\delta_{l,k+m;d}\,\beta^{m}\mathrm{exp}[-\frac{1}{2}(\alpha-\beta)^{2}],$ (4)
where $\delta_{m,k;d}:=\delta_{m,k+d\cdot j}$ with $j\in\mathbb{Z}$. That is, $\delta_{m,k;d}=1$ if and only if $m\equiv k(\mathrm{mod}\ d)$. Also, $\mathcal{N}=\mathcal{N}(\alpha)=\sqrt{\sum_{k=0}^{n}\frac{(n!)^{2}}{k!((n-k)!)^{2}}\alpha^{2(n-k)}}$ is a normalization factor from photon additions. The normalization factor $\mathcal{N}\approx\alpha^{n}$ in the asymptotic limit of $\alpha\gg 1$. In such a limit the optimal choice of $\beta$ yielding maximal overlap becomes $\alpha$, and in that case, fidelity is strictly 1. Even if we consider a more realistic regime with smaller amplitudes($\alpha<9$), we can achieve $F>0.99$ with well-chosen $\beta$ up to $\approx$12 successive photon additions.
@@ II.2 Superposition of coherent states
As the name suggests, the superposition of coherent states (SCS) is a linear combination of coherent states with distinct phases. Typical SCS qubit basis, also known as even/odd cat states are given by
@@
It is worth noting that individual coherent components are not orthogonal, making the normalization factor of SCS depend on the coherent state amplitude $\alpha$. Moreover, orthogonality between |$\ket{\mathcal{C}^{0}_{\alpha}}$ and |$\ket{\mathcal{C}^{1}_{\alpha}}$ is strictly guaranteed by relative phase between their components regardless of their nonorthogonality. This can be explicitly proven by attaining the number basis expansion of 2-dimensional SCSs.
@@
$N_{k,\alpha,d}=\frac{1}{\sqrt{d\sum_{n=0}^{d-1}\omega_{d}^{-kn}\mathrm{exp}[-\alpha^{2}(1-\omega_{d}^{n})]}}.$ (8)
@@
$\begin{split}\bra{m}\ket{\mathcal{C}^{k}_{\alpha,d}}&=N_{k,\alpha,d}\sum_{n}\omega_{d}^{-kn}\bra{m}\ket{\alpha\omega_{d}^{n}}\\ &=N_{k,\alpha,d}\sum_{n}\omega_{d}^{-kn}\frac{(\alpha\omega_{d}^{n})^{m}}{\sqrt{m!}}\\ &=\frac{\alpha^{m}N_{k,\alpha,d}}{\sqrt{m!}}\sum_{n}\omega_{d}^{n(m-k)}=\frac{\alpha^{m}N_{k,\alpha,d}}{\sqrt{m!}}\delta_{m,k;d}.\end{split}$ (9)
@@
III Amplification of Hybrid Entangled states
In this section, we compare the two amplification schemes applied to HESs. The first is addition-subtraction, namely $\hat{a}\hat{a}^{\dagger}$ scheme, while the second comprises successive photon addition, namely $a^{\dagger}{}^{2}$ scheme. Deterministic implementations of both schemes are fundamentally forbidden since they are non-unitary. Thus, quantum channels enabling the probabilistic application of those operations are designed to realize those schemes.
@@ III Amplification of Hybrid Entangled states
|$\begin{split}&\ket{\mathcal{H}^{k,\hat{a}\hat{a}^{\dagger}}_{\alpha,d}}=N_{HES}^{\hat{a}\hat{a}^{\dagger}}(\alpha,d)\,(I\otimes\hat{a}{\hat{a}^{\dagger}})\frac{1}{\sqrt{d}}\sum_{n}\omega_{d}^{-kn}\ket{n,\alpha\omega_{d}^{n}},\\ &\ket{\mathcal{H}^{k,(\hat{a}^{\dagger})^{2}}_{\alpha,d}}=N_{HES}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,d)\,(I\otimes(\hat{a}^{\dagger})^{2})\frac{1}{\sqrt{d}}\sum_{n}\omega_{d}^{-kn}\ket{n,\alpha\omega_{d}^{n}},\end{split}$ (10)
where $N_{HES}^{\hat{a}\hat{a}^{\dagger}}(\alpha,d),N_{HES}^{(\hat{a}^{\dagger})^{2}}(\alpha,d)$ normalization factors required for the consequent states to be normalized. Note that we only operate on CV parts.
$N_{HES}^{\hat{a}\hat{a}^{\dagger}}(\alpha,d)=\frac{1}{\sqrt{\alpha^{4}+3\alpha^{2}+1}},$ (11)
$N_{HES}^{(\hat{a}^{\dagger})^{2}}(\alpha,d)=\frac{1}{\sqrt{\alpha^{4}+4\alpha^{2}+2}}.$ (12)
@@
$\langle P(\hat{a},\hat{a}^{\dagger})\rangle_{\mathcal{H}^{k}_{\alpha,d}}=\langle P(\hat{a},\hat{a}^{\dagger})\rangle_{\alpha}.$ (13)
The trace preserving property of quantum channels and Proposition 1 give an expression for normalization factor $N_{HES}^{P(\hat{a},{\hat{a}^{\dagger}})}(\alpha,d)$:
$N_{HES}^{P(\hat{a},{\hat{a}^{\dagger}})}(\alpha,d)=\frac{1}{\sqrt{\bra{\alpha}(P(\hat{a},{\hat{a}^{\dagger}}))^{\dagger}P(\hat{a},{\hat{a}^{\dagger}})\ket{\alpha}}},$ (14)
@@
[…]
$F^{\hat{a}\hat{a}^{\dagger}}_{HES}=\absolutevalue{\innerproduct{\mathcal{H}^{k}_{g\alpha,d}}{\mathcal{H}^{k,\hat{a}\hat{a}^{\dagger}}_{\alpha,d}}}^{2},$ (15)
$F^{\hat{a}^{\dagger}{}^{2}}_{HES}=\absolutevalue{\innerproduct{\mathcal{H}^{k+2}_{g\alpha,d}}{\mathcal{H}^{k,(\hat{a}^{\dagger})^{2}}_{\alpha,d}}}^{2},$
where $k+2$ for the target state of $F^{\hat{a}^{\dagger}{}^{2}}_{HES}$ is chosen to optimize the fidelity for the schme of two-photon addition.
The optimized gain, G, can be obtained by optimizing $g$ for the coherent amplitude achieving the highest fidelity. For the $\hat{a}\hat{a}^{\dagger}$ scheme, the fidelity is analytically obtained with respect to an arbitrary value of gain $g$ as
$F^{\hat{a}\hat{a}^{\dagger}}_{HES}=\frac{g^{2}\alpha^{4}+2g\alpha^{2}+1}{\alpha^{4}+3\alpha^{2}+1}\mathrm{exp}[-\alpha^{2}(g-1)^{2}],$ (16)
@@
$F^{\hat{a}^{\dagger}{}^{2}}_{HES}=\frac{g^{4}\alpha^{4}}{\alpha^{4}+4\alpha^{2}+2}\mathrm{exp}[-\alpha^{2}(g-1)^{2}].$ (17)
Figure 1: Several measures for the amplified HESs against input coherent amplitudes, $\alpha$, for comparison. (a) Fidelities F of the two schemes: $F_{HES}^{\hat{a}\hat{a}^{\dagger}}$(blue, solid) and $F_{HES}^{\hat{a}^{\dagger}{}^{2}}$(orange, dashed). (b) Optimized gains G of the two schemes: $G_{HES}^{\hat{a}\hat{a}^{\dagger}}$(solid blue) and $G_{HES}^{\hat{a}^{\dagger}{}^{2}}$(orange dashed). (c) Normalized quantum Fisher information, $\mathcal{F}_{N}$, of amplified HESs with respect to input HESs: $\mathcal{F}_{HES}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{HES}$ (solid blue) and $\mathcal{F}_{HES}^{\hat{a}^{\dagger}{}^{2}}/\mathcal{F}_{HES}$(orange dashed). (d) Ratio of quantum Fisher information of the two schemes, $\mathcal{R}=\mathcal{F}_{HES}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{HES}^{\hat{a}^{\dagger}{}^{2}}$.
The fidelities of the two schemes are plotted in Fig. 1(a). One can see that the $\hat{a}\hat{a}^{\dagger}$ scheme yields higher fidelity than $(\hat{a}^{\dagger})^{2}$ scheme. We first note that the fidelities in Eqs. (16) and (17) are the same with the fidelities between input and amplified coherent states due to Proposition 1. An explicit calculation of fidelities in Appendix A indeed has the form of Eq. (13). The behaviors of fidelities between the target coherent state with greater amplitude and the coherent state amplified with either the $\hat{a}\hat{a}^{\dagger}$ scheme or $\hat{a}^{\dagger}{}^{2}$ scheme are known Park16 .
@@
$\langle\mathcal{H}_{\beta,d}^{k+l}|Q(\hat{a},\hat{a}^{\dagger};l)|\mathcal{H}_{\alpha,d}^{k}\rangle=\langle\beta|Q(\hat{a},\hat{a}^{\dagger};l)|\alpha\rangle.$ (19)
Unlike Proposition 1, this agreement between the fidelity of the HES amplification and that of the coherent-state amplification does not solely originate from the orthogonality of discrete variables. We both need discrete variables |$\ket{n}$ which act as labels of continuous variables |$\ket{\alpha\omega_{d}^{n}}$ and equal phase factor $\omega$ which constitutes both qudit number $k$ by relative phase $\omega^{-kn}$ and phase of coherent states |$\ket{\alpha\omega^{n}}$ to cancel out relative phase.
We now obtain the optimized gain defined as the ratio between the amplitudes of the target and initial states giving the maximum fidelity. This can be obtained by solving the equation